OpenStudy (anonymous):

how do you find the midpoint with two ordered pairs, (12,-11), (5,13)

OpenStudy (anonymous):

Just think, first, about what a middle point between an x value of 12 and an x value of 5 might be. Don't worry about the y for now... what is the x value of the middle point between those two x values?

OpenStudy (anonymous):

so its 12+5/2?....

OpenStudy (anonymous):

(12 + 5)/2, yes... I think that's what you meant. Can you find the mid point between the y values?

OpenStudy (anonymous):

so for x its 8.5 and for y its 1 right?

OpenStudy (anonymous):

Looks good to me... always smart to think about where that mid point would actually be in relation to the two endpoints... if you accidentally made a math mistake, your midpoint would be somewhere not between the other 2 points. a quick sketch should tell you if it seems right.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can i ask you another question?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

\[(-1,1), (-1+3\sqrt{3}, 4)\) how do i combine the -1 and the 3\[\sqrt{3}\]

OpenStudy (anonymous):

sry how to i combine the square root?

OpenStudy (anonymous):

you add: -1 + (-1 + 3sqrt(3) ) = -1 - 1 + 3sqrt(3) = -2 + 3sqrt(3) then divide the whole thing by 2: (-2 + 3sqrt(3) ) / 2 leaving you: -2/2 + (3/2)sqrt(3)

OpenStudy (anonymous):

you don't combine the square root... For example, it's sort of like saying "how do I divide x + 2 by 2?" You don't combine the x and the 2... they are already combined by the + symbol... you just set it up as (x+2)/2 and if you want, you can simplify to (x/2) + (2/2) = (x/2) + 1